A convex lens of diameter 8.0 cm is used to focus a parallel beam of light of wavelength 620 nm. If the light be focused at a distance of 20 cm from the lens, what would be the radius of the central bright spot formed?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. . For first minimum diffraction, b sin θ = λ Here, θ = 30°, b = 5 cm ⇒ λ = 5 * sin 30° = 5/2 = 2.5 cm.